Bewildered (bwd), funky ear (fe) and dreamy eyes (de) are recessive alleles on different chromosomes. A pure breeding bewildered, dreamy eyed mouse is crossed to a pure breeding funky ear, mouse, and the F1 mice are inbred to produce F2 progeny. What proportion of the F2 progeny are expected to be bewildered, but not funky ear or dreamy eyed?

The answer is 0.141, but can you show me HOW to solve this?
Answer #1

Answer:

bwd / bwd ; fe+/fe+ ; de/de X bwd+/bwd+ ; fe/fe ; de+/de+ ---Parents

bwd+/bwd ; fe+/fe ; de+/de ---------------------------------F1 (wild type)

F1 x F1---Inbreeding

bwd+/bwd x bwd+/bwd = bwd+_ (wild type-3/4) & bwd/bwd ( Bewildered -1/4)

fe+/fe x fe+/fe = fe+_ (wild type -3/4) & (funky ear -1/4)

de+/de x de+/de = de+_ (wild type -3/4) & de/de (dreamy eyes-1/4)

The proportion of the F2 progeny are expected to be bewildered, but not funky ear or dreamy eyed =(bwd/bwd; fe+_ ; de+_ =1/4 * 3/4 * 3/4 = 9/64 = 0.141